let f(x)=2x^2-x-2, g(x)=x+2 Find f[g(2)]
Answer by stanbon(75887)

let f(x)=2x^2-x-2, g(x)=x+2 Find f[g(2)]
--------------
f[g(2)] = f[(2+2)] = f[4] = 2*4^2-4-2 = 32-4-2 = 26
------------------------
Cheers,
Stan H.
========================